Journal of the Association for Computing Machinery, Vol. 43, No. 6, November 1996, pp. 973–1001.

Efficient Routing in Optical Networks

ALOK AGGARWAL

AMOTZ BAR-NOY

DON COPPERSMITH

RAJIV RAMASWAMI

BARUCH SCHIEBER

AND

MADHU SUDAN

Abstract. This paper studies the problem of dedicating routes to connections in optical networks. In optical networks, the vast bandwidth available in an optical fiber is utilized by partitioning it into several channels, each at a different optical wavelength. A connection between two nodes is assigned a specific wavelength, with the constraint that no two connections sharing a link in the network can be assigned the same wavelength. This paper considers optical networks with and without switches, and different types of routing in these networks. It presents optimal or near-optimal constructions of optical networks in these cases and algorithms for routing connections, specifically permutation routing for the networks constructed here.

Categories and Subject Descriptors. F.2.2 [Analysis of Algorithms and Problem Complexity]: Nonnumerical Algorithms and Problems; G.2.1 [Discrete Mathematics]: Combinatorics; G.2.2 [Discrete Mathematics]: Graph Theory

General Terms. Algorithms

Additional Key words and Phrases. Optical networks, routing, wavelength assignment

Copyright 1996 by ACM, Inc. Permission to make digital or hard copies of part or all of this work for personal or classroom use is granted without fee provided that copies are not made or distributed for profit or direct commercial advantage and that copies bear this notice and the full citation on the first page. Copyrights for components of this work owned by others than ACM must be honored. Abstracting with credit is permitted. To copy otherwise, to republish, to post on servers, or to redistribute to lists, requires prior specific permission and/or a fee. Request permissions from Publications Dept, ACM Inc., fax +1 (212) 869-0481, or permissions@acm.org.